Brain Freeze can't remember. Given functions F and G dtermine the domain of f+g: f(x)=5x-8,g(x)=5/x-2
Thank you for you help

Answer by stanbon(75887)

dtermine the domain of f+g: f(x)=5x-8,g(x)=5/x-2
---------------
f+g = 5x-8 + 5/(x-2)
Domain : all Real Numbers except x=2.
============
Cheers,
Stan H.